# Fine-grained Hallucination Detection and Editing for Language Models

### v3

@@ Acknowledgement
We thank Sewon Min for fruitful discussions in the early stages. Akari Asai is supported by the IBM Fellowship. We thank Stability AI for providing computing to train and evaluate the LMs in this work, and Microsoft Accelerate Foundation Models Research Program for the access to Azure and OpenAI APIs. We thank our annotators who provide high-quality annotations for this work. This work was funded in part by the DARPA MCS program through NIWC Pacific (N66001-19-2-4031), NSF IIS-2044660, and gifts from AI2.
@@ Appendix A Details of Human Annotations
Dataset
Annotated Passage Example
WebNLG
A Long Long Way is a novel written by Sebastian Barry. It was first published in 2005 by $<$entity$>$Faber and Faber viking press$<$/entity$>$ , a $<$entity$>$ British American$<$ /entity$>$ publishing company. The novel tells the story of Willie Dunne, an Irish soldier who has a strong sense of loyalty to his country, but also feels torn between his Irishness and his loyalty to the British army during World War I. The novel was critically acclaimed and received several literary awards, $<$contradictory$>$ including the 2005 Dublin Literary Award $<$/contradictory$>$ and was shortlisted for the Man Booker Prize in the same year.
OpenAssistant
$<$subjective$>$ The most luxurious airline flying within the continental United States is generally regarded as JetBlue Airways$<$/subjective$>$ , particularly for its Mint business class service. Mint offers a private suite experience, featuring fully lie-flat seats, ample legroom and storage, premium food and beverage options, access to onboard WiFi and live TV streaming, and turn-down service complete with duvets and pillows.
Instruction Following
Some of Afghanistan’s secondary cricket home grounds include:
@@ Appendix B More Details of Synthetic Data Creation and Training
Genre Topic
Output
Article Community Reinvestment Act
@@
The Community Reinvestment Act (CRA) is a federal law that encourages commercial banks and savings associations to help meet the needs of borrowers in low- and moderate-income neighborhoods. The Act was passed in 1977 by Congress to reduce discriminatory credit practices against low-income neighborhoods.
Essay National Dodgeball League
@@
The National Dodgeball League is the only professional dodgeball league in the US, and was founded in 2004. It is headed by Commissioner Edward Prentiss out of Hopkins, MN and is composed of 24 professional teams that are divided into the National and American Dodgeball Conferences. In the past, the tournament has featured teams from Canada, Denmark, UK, Japan, New Zealand and Australia.
Blog Post Red Channels
@@
In the beginning of the 1950s, the ”Counterattack” journal published a document, ”Red Channels,” which named 151 individuals in the entertainment industry who were believed to be under Communist manipulation or association. Many of these individuals were being denied employment or had their careers negatively impacted due to their political beliefs or history. ”Red Channels” effectively created a blacklist of those named in the document, further limiting their opportunities in the industry.
Social Media Post Sandman
@@
Check out The Sandman - a semi-retired American professional wrestler! He is best known for his career with Extreme Championship Wrestling (ECW), where he held the ECW World Heavyweight Championship a record five times. He also had stints in World Championship Wrestling (WCW), Total Nonstop Action Wrestling (TNA), and World Wrestling Entertainment (WWE). #Wrestling #ECW #WCW #TNA #WWE
News Report I Don’t Mind (James Brown song)
@@
James Brown’s ”I Don’t Mind” has made its way up the R&B ”Billboard” charts, hitting number four. The track was released in 1961 and also reached number 47 on the Pop ”Billboard” charts. Fans will also be able to hear the song on Brown and the Famous Flames album Live at the Apollo.
Summary Robert Lui
@@
Robert Lui was born on February 23, 1990, and is a professional rugby league footballer from Australia. He plays either as a halfback or five-eighth for the Townsville Blackhawks in the Queensland Cup.
Speech George Sperling
@@
Dear fellow Americans, I am here to propose a solution to improve American Sign Language communication. George Sperling suggests that with a sevenfold reduction in the bandwidth for video transmission, we can achieve this. He even argued that the telephone was originally created for the hearing impaired but it became popularized by the hearing community. Let us not forget our roots and make a change for the better.
Presentation Intro Blissful Ignorance Effect
@@
Have you ever wondered why sometimes people who know less about a product seem to enjoy it more than those who have researched it thoroughly? This phenomenon is called the Blissful Ignorance Effect and it’s a fascinating topic in consumer behavior studies. Our presentation today will explore this effect and why it happens.
Brochure Nidulariaceae Fungi
@@
Explore the fascinating world of Nidulariaceae fungi! This family, found in most ecological regions, includes five different genera, each with its own unique characteristics. With their tiny egg-filled structures, these fungi are a wonder to behold!
Text Message Ashes and Diamonds
@@
Just found out about this book called Ashes and Diamonds by Jerzy Andrzejewski. It’s set during the last few days of WWII. The main character, Maciek, has to kill a Communist soldier. Sounds intense!
@@
Our base model is Llama2-Chat 7b trained using 4xA40 GPUs. Our training code is based off Open-Instruct Wang et al. (2023a) [^13] https://github.com/allenai/open-instruct/blob/main/scripts/finetune_with_accelerate.sh. Table 9 shows the training hyperparameters.
@@ C.1 Detection Results
Table  10 reports the detection results for the Llama2-Chat 7B generations on our curated datasets.
@@
Tables 11 and 12 report the precision and recall on our curated datasets.
@@ C.2 Benchmark Prompts
Figure C.2 shows the prompt used for baseline models when evaluating on the detection benchmark.
@@ C.3 Details of Human Evaluation on Model Outputs
We conduct a human evaluation on 50 detection results of the top two models, namely Fava, and Rt-ChatGPT. Our results are shown in Table  14. We anonymize the model results to avoid potential biases during evaluations. Annotators count how many of the detection and editing results are correct given extensive web searches.
@@ C.4 Manual Analysis on Generated Data for Fava training
We conduct human evaluations on 50 generated data to assess the automatic data creation quality. Prior work introduces entity-centric synthetic data creation Longpre et al. (2021) and often results in unrealistic edits that powerful LMs can easily identify which are synthetically perturbed. Therefore, we evaluate not only validity but also quality. In particular, we evaluate a generated instance from two aspects: (1) validity—whether the model edits violate our annotation scheme; (2) quality—the inserted errors are feasible and realistic and notably different from provided few-shot samples. We ask human annotators to score each category for each passage either 0, 1, or 2 (higher is better; see the detailed criteria in our Appendix).
@@
Table  15 shows examples of generated erroneous data from our synthetic data creation pipeline. We ask human annotators to score each category for each passage either 0, 1, or 2. These scores were defined as follows: a score of 0 indicated that less than half of the inserted errors aligned with the given criteria, a score of 1 signified that at least half of the errors met the established criteria, and a score of 2 indicated that all inserted errors met the criteria. As a comparison, we also generate training data with one-shot generation.
Original Passage
Automatic Data Generations for Training
“Solvalou,” the hyper-entertainment machine known for its sit-down cabinet design and 3D shooting gameplay, has gained critical acclaim. Although it failed to find commercial success, the game recently made a comeback with a digital re-release for the Japanese Wii Virtual Console in 2009 as one of the first titles under the Virtual Console Arcade brand. Despite an announced home port for the 3DO Interactive Multiplayer in 1994, the release was ultimately canceled and never reached the market.
“Solvalou,” the hyper-entertainment machine known for its sit-down cabinet design and 3D shooting gameplay, has gained critical acclaim. ¡invented$>$ The game won the Shooting Star Award in 2020 for its innovative concept.$<$/invented$>$ Although it failed to find commercial success, the game recently made a comeback with a digital re-release for the Japanese Wii Virtual Console in 2009 as one of the first titles under the Virtual Console Arcade brand. Despite an announced home port for the 3DO Interactive $<$entity$>$ $<$delete$>$ Multiplayer$<$/delete$>$ $<$mark$>$ Singleplayer $<$/mark$>$ $<$/entity$>$ in 1994, the release was ultimately canceled and never reached the market. $<$subjective$>$ The game is a masterpiece that revolutionized the arcade industry and will always be remembered as a classic.$<$/subjective$>$
The knapsack problem is a problem in combinatorial optimization where the goal is to determine the optimal selection of items to maximize the total value within a given weight constraint.
The knapsack problem is a problem in $<$entity$>$ $<$delete$>$ combinatorial optimization$<$/delete$>$ $<$mark$>$ sequential search$<$/mark$>$ $<$/entity$>$ where the goal is to determine the optimal selection of items to maximize the total $<$entity$>$ ¡delete$>$ value¡/delete$>$ ¡mark$>$ weight¡/mark$>$ ¡/entity$>$ within a given weight constraint.

### v4

@@ Acknowledgement
We thank Sewon Min for fruitful discussions in the early stages and UW NLP members and annonymous reviewers for their insightful feedback on the draft. We thank Stability AI for providing computing to train and evaluate the LMs in this work, and Microsoft Accelerate Foundation Models Research Program for the access to Azure and OpenAI APIs. We thank our annotators who provide high-quality annotations for this work. This work was funded in part by the DARPA MCS program through NIWC Pacific (N66001-19-2-4031), NSF IIS-2044660, and gifts from AI2. We gratefully acknowledge support from NSF CAREER Grant No. IIS2142739, NSF Grants No. IIS2125201, IIS2203097, and gift funding from Google, MSR, and OpenAI.
@@ Appendix A Details of Human Annotations
Dataset Annotated Passage Example
WebNLG A Long Long Way is a novel written by Sebastian Barry. It was first published in 2005 by $<$entity$>$Faber and Faber viking press$<$/entity$>$ , a $<$entity$>$ British American$<$ /entity$>$ publishing company. The novel tells the story of Willie Dunne, an Irish soldier who has a strong sense of loyalty to his country, but also feels torn between his Irishness and his loyalty to the British army during World War I. The novel was critically acclaimed and received several literary awards, $<$contradictory$>$ including the 2005 Dublin Literary Award $<$/contradictory$>$ and was shortlisted for the Man Booker Prize in the same year.
OpenAssistant $<$subjective$>$ The most luxurious airline flying within the continental United States is generally regarded as JetBlue Airways$<$/subjective$>$ , particularly for its Mint business class service. Mint offers a private suite experience, featuring fully lie-flat seats, ample legroom and storage, premium food and beverage options, access to onboard WiFi and live TV streaming, and turn-down service complete with duvets and pillows.
Instruction Following Some of Afghanistan’s secondary cricket home grounds include:
@@ Appendix B More Details of Synthetic Data Creation and Training
Quality of error insertions.
In our pilot studies, we found that asking ChatGPT to insert multiple error types at the same time easily makes the model misunderstand or get confused with different error types (e.g., swapping or incorrectly identifying an error type). We also found while ChatGPT is capable of generating more simple types of errors such as  Entity , it struggles with generating plausible and difficult perturbations for more nuanced types.
@@
Genre Topic Output
@@
Article Community Reinvestment Act The Community Reinvestment Act (CRA) is a federal law that encourages commercial banks and savings associations to help meet the needs of borrowers in low- and moderate-income neighborhoods. The Act was passed in 1977 by Congress to reduce discriminatory credit practices against low-income neighborhoods.
@@
Essay National Dodgeball League The National Dodgeball League is the only professional dodgeball league in the US, and was founded in 2004. It is headed by Commissioner Edward Prentiss out of Hopkins, MN and is composed of 24 professional teams that are divided into the National and American Dodgeball Conferences. In the past, the tournament has featured teams from Canada, Denmark, UK, Japan, New Zealand and Australia.
@@
Blog Post Red Channels In the beginning of the 1950s, the ”Counterattack” journal published a document, ”Red Channels,” which named 151 individuals in the entertainment industry who were believed to be under Communist manipulation or association. Many of these individuals were being denied employment or had their careers negatively impacted due to their political beliefs or history. ”Red Channels” effectively created a blacklist of those named in the document, further limiting their opportunities in the industry.
@@
Social Media Post Sandman Check out The Sandman - a semi-retired American professional wrestler! He is best known for his career with Extreme Championship Wrestling (ECW), where he held the ECW World Heavyweight Championship a record five times. He also had stints in World Championship Wrestling (WCW), Total Nonstop Action Wrestling (TNA), and World Wrestling Entertainment (WWE). #Wrestling #ECW #WCW #TNA #WWE
@@
News Report I Don’t Mind (James Brown song) James Brown’s ”I Don’t Mind” has made its way up the R&B ”Billboard” charts, hitting number four. The track was released in 1961 and also reached number 47 on the Pop ”Billboard” charts. Fans will also be able to hear the song on Brown and the Famous Flames album Live at the Apollo.
@@
Summary Robert Lui Robert Lui was born on February 23, 1990, and is a professional rugby league footballer from Australia. He plays either as a halfback or five-eighth for the Townsville Blackhawks in the Queensland Cup.
@@
Speech George Sperling Dear fellow Americans, I am here to propose a solution to improve American Sign Language communication. George Sperling suggests that with a sevenfold reduction in the bandwidth for video transmission, we can achieve this. He even argued that the telephone was originally created for the hearing impaired but it became popularized by the hearing community. Let us not forget our roots and make a change for the better.
@@
Presentation Intro Blissful Ignorance Effect Have you ever wondered why sometimes people who know less about a product seem to enjoy it more than those who have researched it thoroughly? This phenomenon is called the Blissful Ignorance Effect and it’s a fascinating topic in consumer behavior studies. Our presentation today will explore this effect and why it happens.
@@
Brochure Nidulariaceae Fungi Explore the fascinating world of Nidulariaceae fungi! This family, found in most ecological regions, includes five different genera, each with its own unique characteristics. With their tiny egg-filled structures, these fungi are a wonder to behold!
@@
Text Message Ashes and Diamonds Just found out about this book called Ashes and Diamonds by Jerzy Andrzejewski. It’s set during the last few days of WWII. The main character, Maciek, has to kill a Communist soldier. Sounds intense!
@@
Our base model is Llama2-Chat 7b trained using 4xA40 GPUs. Our training code is based off Open-Instruct (Wang et al., 2023a) [^15] https://github.com/allenai/open-instruct/blob/main/scripts/finetune_with_accelerate.sh. Table 11 shows the training hyperparameters.
@@ C.1 Detection Results
Table  12 reports the detection results for the Llama2-Chat 7B generations on our curated datasets.
@@
Tables 13 and 14 report the precision and recall on our curated datasets.
@@ C.2 Benchmark Prompts
Table 15 shows the prompt used for baseline models when evaluating on the detection benchmark.
@@ C.3 Details of Human Evaluation on Model Outputs
We conduct a human evaluation on 50 detection results of the top two models, namely Fava, and Rt-ChatGPT. Our results are shown in Table  16. We anonymize the model results to avoid potential biases during evaluations. Annotators count how many of the detection and editing results are correct given extensive web searches.
@@ C.4 Manual Analysis on Generated Data for Fava training
We conduct human evaluations on 50 generated data to assess the automatic data creation quality. Prior work introduces entity-centric synthetic data creation Longpre et al. (2021) and often results in unrealistic edits that powerful LMs can easily identify which are synthetically perturbed. Therefore, we evaluate not only validity but also quality. In particular, we evaluate a generated instance from two aspects: (1) validity—whether the model edits violate our annotation scheme; (2) quality—the inserted errors are feasible and realistic and notably different from provided few-shot samples. We ask human annotators to score each category for each passage either 0, 1, or 2 (higher is better).
@@
Table  17 shows examples of generated erroneous data from our synthetic data creation pipeline. We ask human annotators to score each category for each passage either 0, 1, or 2. These scores were defined as follows: a score of 0 indicated that less than half of the inserted errors aligned with the given criteria, a score of 1 signified that at least half of the errors met the established criteria, and a score of 2 indicated that all inserted errors met the criteria. As a comparison, we also generate training data with one-shot generation.
Original Passage Automatic Data Generations for Training
“Solvalou,” the hyper-entertainment machine known for its sit-down cabinet design and 3D shooting gameplay, has gained critical acclaim. Although it failed to find commercial success, the game recently made a comeback with a digital re-release for the Japanese Wii Virtual Console in 2009 as one of the first titles under the Virtual Console Arcade brand. Despite an announced home port for the 3DO Interactive Multiplayer in 1994, the release was ultimately canceled and never reached the market. “Solvalou,” the hyper-entertainment machine known for its sit-down cabinet design and 3D shooting gameplay, has gained critical acclaim. ¡invented$>$ The game won the Shooting Star Award in 2020 for its innovative concept.$<$/invented$>$ Although it failed to find commercial success, the game recently made a comeback with a digital re-release for the Japanese Wii Virtual Console in 2009 as one of the first titles under the Virtual Console Arcade brand. Despite an announced home port for the 3DO Interactive $<$entity$>$ $<$delete$>$ Multiplayer$<$/delete$>$ $<$mark$>$ Singleplayer $<$/mark$>$ $<$/entity$>$ in 1994, the release was ultimately canceled and never reached the market. $<$subjective$>$ The game is a masterpiece that revolutionized the arcade industry and will always be remembered as a classic.$<$/subjective$>$
The knapsack problem is a problem in combinatorial optimization where the goal is to determine the optimal selection of items to maximize the total value within a given weight constraint. The knapsack problem is a problem in $<$entity$>$ $<$delete$>$ combinatorial optimization$<$/delete$>$ $<$mark$>$ sequential search$<$/mark$>$ $<$/entity$>$ where the goal is to determine the optimal selection of items to maximize the total $<$entity$>$ $<$delete$>$ value$<$/delete$>$ $<$mark$>$ weight$<$/mark$>$ $<$/entity$>$ within a given weight constraint.
@@ C.5 More Ablations
Analysis on data scaling.
We conduct ablations of different data scaling in Fava. Figure 7 shows the scaling results.
Figure 7: Training Size